a hiker hikes up a mountain 1 mi/h slower than she hikes down the mountain. if it takes her 1.5 hours to hike up the mountain and only 1 hour to hike down it, how fast does she move in each direction?
Answer by mananth(16949)

a hiker hikes up a mountain 1 mi/h slower than she hikes down the mountain. if it takes her 1.5 hours to hike up the mountain and only 1 hour to hike down it, how fast does she move in each direction?
hike down speed = x mph
hike up speed = x-1 mph
...
distance is same up & down
...
d= rt
...
(x-1)*1.5 = 1*x
1.5x-1.5 = x
1.5x-x=1.5
0.5x=1.5
/0.5
x=3 mph rate hiking down
hiking up rate = x-1 = 3-1 = 2 mph
...
m.ananth@hotmail.ca